OpenStudy (sepeario):

How many 3-digit numbers are there where the sum of the digits equals 11?

OpenStudy (kainui):

Any ideas? Can you come up with maybe a specific answer that works? This seems to be closely related to how many ways can you arrange 3 numbers.

OpenStudy (sepeario):

I wrote out all the combinations (1+1+9, 1+2+8, etc.) which therefore gives 9 possibilites, and multiplied it by 6, because you can invert a 3 digit number 6 ways.

ganeshie8 (ganeshie8):

are you familiar with stars and bars ?

OpenStudy (sepeario):

no @ganeshie8

ganeshie8 (ganeshie8):

stars and bars directly gives you the answer, but lets work it by counting : say the 3 digit number is \(abc\), such that : \[a+b+c = 11 \\ 0\lt a \le 9\\0\le b\le 9\\ 0\le c \le 9\]

OpenStudy (sepeario):

Ok.

OpenStudy (sepeario):

Honestly I don't know what that means at all, sorry

OpenStudy (sepeario):

@ganeshie8

ganeshie8 (ganeshie8):

lets say a = 1, then we get : \[\large b+c = 10\] can you guess how many ways you can cook up b and c ?

OpenStudy (sepeario):

isn't it 5?

ganeshie8 (ganeshie8):

whats the minimum possible value of \(b \) ?

OpenStudy (sepeario):

1

ganeshie8 (ganeshie8):

for sure, \(b\) cannot be 0 (why ?) \(b\) has to range from 1 to 9 and for each of this value of b, c us uniquely determined. so total number of possible ways for b+c=10 is 9

OpenStudy (sepeario):

oh yeah because you can swap them.

ganeshie8 (ganeshie8):

Exactly ! corrected some typoes : for sure, \(b\) cannot be \(0\) (why ?) \(b\) has to range from \(1\) to \(9\) and for each of this value of \(b\), the value of \(c\) can be uniquely determined. so total number of possible ways for \(b+c=10\) is \(9\)

OpenStudy (sepeario):

yep.

ganeshie8 (ganeshie8):

a=1 : you get 9 possible numbers can you work a=2 ?

OpenStudy (sepeario):

8

ganeshie8 (ganeshie8):

hmm lets see... when a = 2 : \[\large b+c = 9\] whats the minimum value of \(b\) ? whats the maximum value of \(b\) ?

OpenStudy (sepeario):

maximum 8 min 1

ganeshie8 (ganeshie8):

think again why can't the value of \(b\) be 0 or 9 ?

OpenStudy (sepeario):

oh yeah!

ganeshie8 (ganeshie8):

a=2, b+c=9 gives you 10 numbers

ganeshie8 (ganeshie8):

we can sit and count all remaining cases (a=3,4,5,6,7,8,9) or look for a pattern...

ganeshie8 (ganeshie8):

lets work a=3 case, then we will see a pattern may be.. .

OpenStudy (sepeario):

ok, there are 7 possibilites.

OpenStudy (sepeario):

actually aren't there nine? because 0 and 8 can be used.

ganeshie8 (ganeshie8):

exactly ! looks we have got a pattern :3

OpenStudy (sepeario):

ok, so for every possibility of a there are nine possibilities for b and c?

ganeshie8 (ganeshie8):

a=1 : b+c=10 ( 9 possible numbers) a=2 : b+c= `9` ( `10` possible numbers) a=3 : b+c= `8` ( `9` possible numbers) a=4 : b+c= `7` ( `8` possible numbers) ....

ganeshie8 (ganeshie8):

^^

OpenStudy (sepeario):

oh ok.

ganeshie8 (ganeshie8):

for a=8, b+c = `3`, you will get `4` possible numbers

ganeshie8 (ganeshie8):

In general : the equation \(b+c = n\) has \(n+1\) possible non negative solutions

OpenStudy (sepeario):

yes, except for the first one.

ganeshie8 (ganeshie8):

yes :) thats because we cannot allow "10" for b or c so whats the final answer ?

OpenStudy (sepeario):

61!

OpenStudy (sepeario):

So how is this related to the equation that you were talking about before?

ganeshie8 (ganeshie8):

thats just a lazy way of expressing the sum : 9 + 10 + 9 + 8 + ... + 3 right ?

OpenStudy (sepeario):

Yes.

OpenStudy (sepeario):

So how do u figure it though?

ganeshie8 (ganeshie8):

above sum be represented as below : \[\large \sum \limits_{a=1}^9 ((11-a) +1) - 2\]

OpenStudy (sepeario):

How would I get that in the first place?

ganeshie8 (ganeshie8):

that comes from solving a simpler equation : \(\large x+y = n\)

ganeshie8 (ganeshie8):

how many solutions exist in non negative integers ?

ganeshie8 (ganeshie8):

the answer is \(\large n+1\), right ?

OpenStudy (sepeario):

yes, but we didn't know that before we figured out the answer.

ganeshie8 (ganeshie8):

thats right ! based on that can u cookup a formula for number of solutions of \(\large x+y \le n\) ?

ganeshie8 (ganeshie8):

above formula is almost same as the question you have originally posted (a+b+c=11)

OpenStudy (sepeario):

i don't understand what the variables represent

ganeshie8 (ganeshie8):

what variables ?

OpenStudy (sepeario):

x, y and n

ganeshie8 (ganeshie8):

okay, let me give a specific value for \(n\) and rephrase the problem : How many "non negative integer solutions" are there for the equation \(\large x+y \le 10\) ?

OpenStudy (sepeario):

ok.

ganeshie8 (ganeshie8):

I'll give you the answer, try the proof in ur free time :) Answer : \(\large \sum \limits_{k=0}^{10} (k+1 )\)

OpenStudy (sepeario):

what does the e mean

ganeshie8 (ganeshie8):

what e ?

OpenStudy (sepeario):

the equation sign

ganeshie8 (ganeshie8):

\[\large \sum \limits_{k=0}^4 k \] is just a lazy way of writing : \[\large 0 + 1+2+3+4\]

ganeshie8 (ganeshie8):

this site explains that SUM notation better : http://www.mathsisfun.com/algebra/sigma-notation.html

OpenStudy (sepeario):

what is he k on the right

OpenStudy (sepeario):

oh ok i understand now.

OpenStudy (sepeario):

so how in the first place did u formulate the equation

ganeshie8 (ganeshie8):

good :)

ganeshie8 (ganeshie8):

Spoiler : proof @ http://openstudy.com/users/ganeshie8#/updates/53cbc2dee4b0bff525b1e8f1

OpenStudy (sepeario):

isn't that just your updates page?

OpenStudy (sepeario):

so i don't understand how you made the equation @ganeshie8

ganeshie8 (ganeshie8):

scroll down a bit to see the actual proof

OpenStudy (sepeario):

ok thank you very much @ganeshie8

ganeshie8 (ganeshie8):

np :) let me know if something is not clear in that page